# Exact mobility edges in quasiperiodic network models with slowly varying potentials

### v2

@@ III.1 Hermitian models
where $d_{i,k}^{\dagger}$ ($d_{i,k}$) is the creation (annihilation) operator at the $k$-th site of cell $i$, $V_{i}$ denotes the quasiperiodic slowly varying potentials defined as $V_{i}=V\cos(\pi\alpha i^{\nu})$, and $U_{k}$ represents constant potentials. Here we set $U_{k}=0$ for simplicity, and the results for nonzero cases are similar. Fig. 2(b) illustrates the multipath quasiperiodic network model with $N=3$. For such a multipath quasiperiodic model, the MEs can also be obtained through the effective Hamiltonian. To see that, we begin with the eigenvalue equations for the periodic sites
$v_{i-1,k}=\frac{-tu_{i-1}-tu_{i}}{E},\quad v_{i,k}=\frac{-tu_{i}-tu_{i+1}}{E},$ (37)
where |$\ket{\psi^{n}}=\sum_{i}(u_{i}^{n}c_{i}^{\dagger}+\sum_{k}v_{i,k}^{n}d_{i,k}^{\dagger})\ket{0}$. Substituting these results into the eigenvalue equation for $u_{i}$ yields
@@
Further, let us rewrite $u_{i}$ as $u_{j}$ to distinguish the index between the effective chain and original chain, we can obtain the effective Hamiltonian
@@ III.2 Non-Hermitian models
The PT symmetry is realized by balancing the gain and loss in the periodic sites. The eigenvalue equations for the periodic sites are
@@ IV Experimental Proposal
Finally, we discuss the possible way to realize these network models. The realization of quasiperiodic network models with various periodic network sites is challenging in cold-atom systems [64, 65, 66], however, it can be easily realized using the coupled single-mode optical waveguides [67, 68, 19, 69]. By controlling the distance between neighboring waveguides and tuning their refraction index, we can modulate the hopping strength and on-site potential [70, 71], enabling the realization of various quasiperiodic network models with slowly varying potentials. To observe the localized-delocalized transition in such systems, we can inject light into a single waveguide and measure the distribution of light intensity after it propagates through a certain distance in the waveguides array [see Fig. 9(a)]. The propagation of light along the waveguide can be described by tight-binding equation
@@
In Fig. 9(b), we plot the fractal dimension of the eigenstates as a function of energy $E$ and potentail strength $V$ for a small system ($L=100$), where we can also see clear MEs. Here we consider quasiperiodic mosaic model with slowly varying potentials when $\kappa=2$. To simulate the experimental results, we plot the evolution of the wave function for various potential strength $V$ in Figs. 9(c) - 9(f). Although the specific evolution of wave function depends strongly on its initial state, the properties of localized state or extended state remain the same. Here, we set the initial state as $\psi(i,0)=\ket{0,\cdots,1_{i=50},\cdots,0}$. For $V=0.5t$, the wave function gradually spreads through the entire system. As the potential strength increases to $V=t$, the wave function propagating to the boundary decreases, since the number of extended states in the system is reduced. When $V=2t$ and $V=3t$, the wave functions become localized around the center of the system, as all states away from the band center are localized, which we leave as an open question to be investigated in the future. Our recent experiments about the reentrant phenomena in quasiperiodic systems can be find in Ref. [63].
@@ V Conclusions
To conclude, in this work we propose a new class of 1D quasiperiodic network models slowly varying potentials and we can exactly determine their MEs. Unlike the previously proposed quasiperiodic network models [28], this model doesn’t exhibit hidden self-duality, since the slowly varying potential is not self-dual. We integrate out the periodic sites to obtain an effective model, where $g(E)V$ represents the effective potential and $f(E)$ denotes the effective eigenenergy. Substituting $f$ and $g$ into $f=\pm(2t^{\kappa}\pm gV)$ will give the concrete MEs, following Das Sarma et. al. [59]. We confirm these results using numerical method. With this method, we can study the exact MEs with slowly varying potentials in some much more complicated models, including non-Hermitian ones. In any way, the exact MEs in models with slowly varying potentials should greatly broaden our understanding of MEs in 1D disordered models.
@@
Now, it is quite possible that we are in a position to reflect on the possible future of quantum simulations. In this work, we show that the rich phases in the quasiperiodic network models can be simulated using the optical waveguides. In actually, there are much more platforms to simulate these models, including optical and acoustic waveguides [75, 33, 76], ultracold atoms, superconducting qubits [77, 78, 79, 80, 81, 82], in which most of the essential physics in these models can be obtained using some small quantum systems. In recent years, these small systems have been used to demonstrate these fundamental physics. Yet, the detailed physics, including the scaling laws and the phase boundaries have not been precisely determined. This is different from that in the community of condensed matter physics, in which the phase transition and the associated scaling exponents have been precisely determined in experiments. The similar accuracy has not yet been achieved in quantum simulation. In the future, with the increasing of experimental techniques, it is quite possible that these simulating platforms can be used to determine the precise phase boundary and the associated critical exponents, making quantum simulations a powerful tool to explore possible new physics. At the current stage, only some intriguing phenomena have been simulated. If this kind of accuracy can be achieved, the physical models may serve as an important platform for exploring Anderson transition in the network models [83], which will help us to resolve the long-standing problems in this physics. Especially, in the presence of many-body interaction [84, 85, 86] and in the higher dimensional models with various potentials [87, 88], these physics cannot be efficiently simulated using classical computers, however, it is quite possible that these new physics can be simulated using these intriguing platforms.
@@ Acknowledgements.
This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0500000), the National Natural Science Foundation of China (Grant No. 12374017, No. 12074362) and the Innovation Program for Quantum Science and Technology (2021ZD0303303, 2021ZD0301200, 2021ZD0301500). A.-M. G. is supported by the NSFC (Grant Nos. 12274466 and 11874428) and the Hunan Provincial Science Fund for Distinguished Young Scholars (Grant No. 2023JJ10058).

### v3

@@ III.1 Hermitian models
[…]
$Ed_{i-1,k}=[d_{i-1,k},H]=-t(c_{i-1}+c_{i}),$ (38)
the eigenvalue equation for the periodic sites, applying the eigenvector into the above equation, which is equivalent to the replacement
$d_{i,k}\rightarrow v_{i,k},\quad c_{i}\rightarrow u_{i},$ (39)
will yields
$v_{i-1,k}=\frac{-tu_{i-1}-tu_{i}}{E},\quad v_{i,k}=\frac{-tu_{i}-tu_{i+1}}{E}.$ (40)
Substituting these results into the eigenvalue equation for $u_{i}$ from the following operator equation
$Ec_{i}=[c_{i},H]=V_{i}c_{i}-t\sum_{k}(d_{i,k}+d_{i-i,k}),$ (41)
will yields
@@
Here, the coefficient $N$ comes from the $N$ periodic sites; see $N=4$ in Fig. 1(b). Further, let us rewrite $u_{i}$ as $u_{j}$ to distinguish the index between the effective chain and original chain, we can obtain the effective Hamiltonian
@@ III.2 Non-Hermitian models
[…]
$Ed_{i-1,k}=[d_{i-1,k},H]=-t(c_{i-1}+c_{i})+U_{k}d_{i-1,k},$ (46)
and the replacement rule in Eq. 39 will yields
@@
This equation is essentially the same as Eq. 42. Obviously, when the complex potentials are not opposite, this equation will become complex, yielding an possible route for non-Hermitian physics [64, 22].
@@ IV Experimental Proposal
Finally, we discuss the possible way to realize these network models. The realization of quasiperiodic network models with various periodic network sites is challenging in cold-atom systems [65, 66, 67], however, it can be easily realized using the coupled single-mode optical waveguides [68, 69, 70, 71, 72]. By controlling the distance between neighboring waveguides and tuning their refraction index, we can modulate the hopping strength and on-site potential [73, 74, 75], enabling the realization of various quasiperiodic network models with slowly varying potentials. To observe the localized-delocalized transition in such systems, we can inject light into a single waveguide and measure the distribution of light intensity after it propagates through a certain distance in the waveguides array [see Fig. 10(a)]. The propagation of light along the waveguide can be described by tight-binding equation
@@
[…]
@@ V Conclusions
To conclude, in this work we propose a new class of 1D quasiperiodic network models slowly varying potentials and we can exactly determine their MEs. Unlike the previously proposed quasiperiodic network models [27], this model doesn’t exhibit hidden self-duality, since the slowly varying potential is not self-dual. We integrate out the periodic sites to obtain an effective model, where $g(E)V$ represents the effective potential and $f(E)$ denotes the effective eigenenergy. Substituting $f$ and $g$ into $f=\pm(2t^{\kappa}\pm gV)$ will give the concrete MEs, following Das Sarma et. al. [59]. We confirm these results using numerical method. With this method, we can study the exact MEs with slowly varying potentials in some much more complicated models, including non-Hermitian ones. Additionally, we also study the connection between topological transition and localization-delocalization transition induced by a complex phase, as discussed in Refs. [64, 22]. In this way, the exact MEs in models with slowly varying potentials should greatly broaden our understanding of MEs in 1D disordered models.
@@
Now, it is quite possible that we are in a position to reflect on the possible future of quantum simulations. In this work, we show that the rich phases in the quasiperiodic network models can be simulated using the optical waveguides. In actually, there are much more platforms to simulate these models, including optical and acoustic waveguides [80, 33, 81], ultracold atoms, superconducting qubits [82, 83, 84, 85, 86, 87], in which most of the essential physics in these models can be obtained using some small quantum systems. In recent years, these small systems have been used to demonstrate these fundamental physics. Yet, the detailed physics, including the scaling laws and the phase boundaries have not been precisely determined. This is different from that in the community of condensed matter physics, in which the phase transition and the associated scaling exponents have been precisely determined in experiments. The similar accuracy has not yet been achieved in quantum simulation. In the future, with the increasing of experimental techniques, it is quite possible that these simulating platforms can be used to determine the precise phase boundary and the associated critical exponents, making quantum simulations a powerful tool to explore possible new physics. At the current stage, only some intriguing phenomena have been simulated. If this kind of accuracy can be achieved, the physical models may serve as an important platform for exploring Anderson transition in the network models [88], which will help us to resolve the long-standing problems in this physics. Especially, in the presence of many-body interaction [89, 90, 91] and in the higher dimensional models with various potentials [92, 93], these physics cannot be efficiently simulated using classical computers, however, it is quite possible that these new physics can be simulated using these intriguing platforms. In the past ten years, we have witnessed the huge progress in quantum simulation with these waveguides, from two coupled waveguides to large networks without thousands of waveguides, and thus in the next decade, We will optimistically to see quantum simulation, as a new means of research, to unveil new physical phenomena.
@@ Acknowledgements.
This work is supported by the Strategic Priority Research Program of the Chinese Academy of Sciences (Grant No. XDB0500000), the National Natural Science Foundation of China (Grant No. 12374017, No. 12074362 and No. U23A2074) and the Innovation Program for Quantum Science and Technology (2021ZD0303303, 2021ZD0301200, 2021ZD0301500). A.-M. G. is supported by the NSFC (Grant Nos. 12274466 and 11874428) and the Hunan Provincial Science Fund for Distinguished Young Scholars (Grant No. 2023JJ10058).